# A case study of gas impacted by black-hole jets with the JWST: outflows, bow shocks, and high excitation of the gas in the galaxy IC5063

### v1

@@ 4.3 The molecular gas mass
We use the Tex map from S(1) and S(2) to determine the mass in each position of the galaxy, keeping in mind that this Tex indeed describes a (near) LTE temperature up to J=5. Combining this temperature with the S(1) luminosity map, we obtain the mass map shown in Fig. 15. The integrated mass in the FOV, M is 1.0107 . An ortho-to-para ratio of 3 is assumed for this calculation, which is reasonable for temperatures 200 K (e.g., Rigopoulou et al. 2002). A more massive reservoir of gas does plausibly exist (Guillard et al., 2012), but its proper mass computation would require similar S(0) observations, as the mass measurement is highly sensitive to the Tex spatial distribution. For comparison, in a similar aperture, the cold mass from ALMA C0(1-0) data is 8.2(0.6)108 , when plugging the CO(1-0) flux (30.62.4 Jy km s-1; Dasyra et al. 2022) and a Galactic CO luminosity to mass factor =4.6 K-1 km-1 s pc-2 (Bolatto et al., 2013) into Equation 3 of Solomon et al. (1997).
Concerning individual regions of interest, the bow-shock shaped structure contains 8.8105 of warm molecular gas, while the mass around the NW radio lobe is similar, 7.8105 . Of the latter mass, 60(14)% is linked to the outflow in the , as indicated by spectral fitting with a double Gaussian. The computation of the mass flow rate of the accelerated gas requires knowledge of the deprojected velocity, the outflow geometry (Lutz et al., 2019; Veilleux et al., 2020), and the exact outflow starting point(s) thus the distance d traveled by the outflow(s). We performed an order of magnitude calculation, assuming that the outflow(s) started locally, upon impact of the radio jet with clouds near the lobe, for d in the range 60-250 pc (i.e., the difference in the spatial offset between the CO(1–0) peak and the CO(4–3) peak in the ALMA data and the difference between the clump c2 and the NW lobe in the present JWST data), and for an indicative velocity of 100 km s-1. Then, the mass flow rate within 1″ from the NW radio lobe, computed as MVd-1,
is in the range 0.11 yr-1. Similarly, within 1″ from the SE radio lobe, the mass is 7.4105 , the outflow fraction is 40(10)%, and the contribution to the total mass flow rate is just a little lower than that of the NW lobe. Still, these numbers are low enough that the mass flow rate is dominated by the CO and HI emitting gas (Morganti et al., 2007, 2015; Dasyra et al., 2016), even if some were to be unaccounted for due to the lack of S(0) data.
@@ 5 Conclusions
We obtained JWST MIRI data of the nucleus of the nearby galaxy IC5063, in which a jet is known to interact with dense gas clouds and drive multi-phase outflows, changing the gas excitation and cloud stability conditions. Our new JWST data provided IFU datacubes of the inner 32kpc2 of the galaxy (in the common FOV of all observations), and revealed the following findings.
@@
Outflows are identified in more than ten discrete locations of IC5063. These include the vicinity of the two radio lobes, the nucleus, and a biconical structure perpendicular to the jet trail, which host previously known outflows. Additionally, new outflows are detected at points near or intermediate to the above, at the outer edges of the EELR, in the interior of the EELR, and in an ionized gas bubble. Overall, outflows and turbulent kinematics are seen in 2/3rds of the common FOV of our observations. The outflow velocities reach and occasionally exceed 1000 km s-1 for the ionized gas. In the bright [S IV] line, within 600 pc of the nucleus, the outflow velocities often exceed the local escape velocity. For the molecular gas, the observed velocities are lower, typically not exceeding the escape velocity.
@@
Appendix A Notes on the instrument’s behaviour in the presence of bright line emission
A.1 Data showing the cruciform artifact
@@
Besides the PSF residuals that are related to a bright continuum source, there is an artifact that can affect any part of the detectors that suffers from the leakage of photons from a bright emission line along the x axis of the raw data. The issue is reported and depicted for the imager in the JWST user documentation webpage111https://jwst-docs.stsci.edu/known-issues-with-jwst-data/miri-known-issues#MIRIKnownIssues-cruciformCruciformartifact. It is the so-called cruciform artifact. It originates from internal scattering of the light within the MIRI detectors, in particular at lower wavelengths, which have a 27% quantum efficiency. Photons that are neither absorbed nor able to escape (in particular for outer PSF regions) are diffracted by the lattice and cause straight-light patterns (Gaspar et al., 2021; Argyriou et al., 2023). This was occasionally clearly seen in our data, as stripes oriented north-south (e.g., in the reshifted [Ne III] and [Ne V] emission shown in Fig. 16). The cruciform was more easily seen for bright emission lines. Bright lines contaminate neighboring detector slices at a constant y-axis pixel. But both because of the detector distortion and also because neighboring detector slices do not correspond to neighboring sky positions, the contamination appears at different wavelengths from the emission line peak wavelength and in various positions in the reconstructed cubes’ FOV. Moreover, because of diffraction, the stray light can either be seen as a stripe, or as an Airy diffraction pattern (see the [Ne VI] emission in Fig. 16). The pipeline, by default, automatically subtracts a (straight line emission) model in channels 1 and 2, in an attempt to remove this stray light. However, because of the detector distortion, signal may be oversubtracted from some regions and undersubtracted in other regions. This leads to regions of images that appear to have cavities (e.g., [Ne III] in Fig. 16) and/or strong absorption (as in all panels of Fig. 17), that can contain emission blobs within the absorption area. This can be confused for real absorption in case there is a bright, warm background emission source such as a jet’s synchrotron in the background. Or it could be confused for a P-Cygni or inverse P-cygni profile. In our case, there are winds (Dasyra et al., 2015), but there is no detected synchrotron emission in the regions where the absorption is seen. (Note that the images and spectra shown in Fig. 17 are continuum subtracted.) Therefore, the absorption would have to be tracing outflows of cold, yet highly ionized gas that is absorbing the emission of hot dust in the background - a rather unusual configuration. However, we report the presence of the artifact for demonstration purposes.
A.2 Data with signal potentially linked to the cruciform artifact
@@
A.3 Nuclear PSF
The image of the continuum for each of the four MIRI channels is shown in Fig. 19. The figure reveals the considerable domination of the nuclear PSF over any other continuum source in the FOV.

### v2

@@ 4.3 The molecular gas mass
We use the Tex map from S(1) and S(2) to determine the mass in each position of the galaxy, keeping in mind that this Tex indeed describes a (near) LTE temperature up to J=5. Combining this temperature with the S(1) luminosity map, we obtain the mass map shown in Fig. 15. The integrated mass in the FOV, M is 1.0107 . An ortho-to-para ratio of 3 is assumed for this calculation, which is reasonable for temperatures 200 K (e.g., Rigopoulou et al. 2002). A more massive reservoir of gas does plausibly exist (Guillard et al., 2012), but its proper mass computation would require similar S(0) observations, as the mass measurement is highly sensitive to the Tex spatial distribution. For comparison, in a similar aperture, the cold mass from ALMA C0(1-0) data is 8.2(0.6)108 , when plugging the CO(1-0) flux (30.62.4 Jy km s-1; Dasyra et al. 2022) and a Galactic CO luminosity to mass factor =4.6 K-1 km-1 s pc-2 (Bolatto et al., 2013) into Equation 3 of Solomon et al. (1997). For the cold gas, the mass in the outflow is too small (2106 ; Dasyra et al. 2016, to indicate that the bulk of the reservoir requires a sub-Galactic .
Concerning individual regions of interest, the bow-shock shaped structure contains 8.8105 of warm molecular gas, while the mass around the NW radio lobe is similar, 7.8105 . Of the latter mass, 60(14)% is linked to the outflow in the , as indicated by spectral fitting with a double Gaussian. The computation of the mass flow rate of the accelerated gas requires knowledge of the deprojected velocity, the outflow geometry (Lutz et al., 2019; Veilleux et al., 2020), and the exact outflow starting point(s) thus the distance d traveled by the outflow(s). We performed an order of magnitude calculation for the NW radio-lobe outflow, assuming that it started locally, upon impact of the jet with clouds at the radio lobe. We used d in the range 60-250 pc (i.e., the difference in the spatial offset between the CO(1–0) peak and the CO(4–3) peak in the ALMA data and the difference between the clump c2 and the NW lobe in the present JWST data), and an indicative velocity of 100 km s-1. Then, the mass flow rate within 1″ from the NW radio lobe, computed as MVd-1, is in the range 0.11 yr-1. These numbers also indicate an indicative outflow travel time, and thus a mininum jet age, of 0.6-2.4 Myrs. Similarly, within 1″ from the SE radio lobe, the mass is 7.4105 , the outflow fraction is 40(10)%, and the contribution to the total mass flow rate is just a little lower than that of the NW lobe. Still, these numbers are low enough that the mass flow rate is dominated by the CO and HI emitting gas (Morganti et al., 2007, 2015; Dasyra et al., 2016), even if some were to be unaccounted for due to the lack of S(0) data.
@@ 5 Conclusions
We obtained JWST MIRI data of the nucleus of the nearby galaxy IC5063, in which a jet is known to interact with dense gas clouds and drive multi-phase outflows, changing the gas excitation. Our new JWST data provided IFU datacubes of the inner 32kpc2 of the galaxy (in the common FOV of all observations), and revealed the following findings.
@@
Outflows are identified in more than ten discrete locations of IC5063. These include the vicinity of the two radio lobes, the nucleus, and a biconical structure perpendicular to the jet trail, which host previously known outflows. Additionally, new outflows are detected at points near or intermediate to the above, at the outer edges of the EELR, in the interior of the EELR, and in an ionized gas bubble. Overall, outflows and turbulent kinematics are seen in the central 3.5 kpc2. The outflow velocities reach and occasionally exceed 1000 km s-1 for the ionized gas. In the bright [S IV] line, within 600 pc of the nucleus, the outflow velocities often exceed the local escape velocity. For the molecular gas, the observed velocities are lower, typically not exceeding the escape velocity.
@@
Appendix A Spectral line and continuum maps
A.1 Spectral line emission
The emission of all detected spectral lines, integrated over all wavelengths showing relevant emission, is shown in 16. All lines show extended emission. Typically, the emission becomes more nucleated with increasing IP.
![[Uncaptioned image]](x10.png)
Figure 16: Continued.
A.2 Continuum emission
The image of the continuum in each of the four MIRI channels, representing the instrument’s PSF, is shown in Fig. 17. The figure reveals the domination of the PSF over any other continuum source in the FOV.
@@ A.2 Continuum emission
Appendix B Notes on the instrument’s behaviour in the presence of bright line emission
B.1 Data showing the cruciform artifact
@@
Besides the PSF residuals that are related to a bright continuum source, there is an artifact that can affect any part of the detectors that suffers from the leakage of photons from a bright emission line along the x axis of the raw data. The issue is reported and depicted for the imager in the JWST user documentation webpage444https://jwst-docs.stsci.edu/known-issues-with-jwst-data/miri-known-issues#MIRIKnownIssues-cruciformCruciformartifact. It is the so-called cruciform artifact. It originates from internal scattering of the light within the MIRI detectors, in particular at lower wavelengths, which have a 27% quantum efficiency. Photons that are neither absorbed nor able to escape (in particular for outer PSF regions) are diffracted by the lattice and cause straight-light patterns (Gaspar et al., 2021; Argyriou et al., 2023). This was occasionally clearly seen in our data, as stripes oriented north-south (e.g., in the reshifted [Ne III] and [Ne V] emission shown in Fig. 18). The cruciform was more easily seen for bright emission lines. Bright lines contaminate neighboring detector slices at a constant y-axis pixel. But both because of the detector distortion and also because neighboring detector slices do not correspond to neighboring sky positions, the contamination appears at different wavelengths from the emission line peak wavelength and in various positions in the reconstructed cubes’ FOV. Moreover, because of diffraction, the stray light can either be seen as a stripe, or as an Airy diffraction pattern (see the [Ne VI] emission in Fig. 18). The pipeline, by default, automatically subtracts a (straight line emission) model in channels 1 and 2, in an attempt to remove this stray light. However, because of the detector distortion, signal may be oversubtracted from some regions and undersubtracted in other regions. This leads to regions of images that appear to have cavities (e.g., [Ne III] in Fig. 18) and/or strong absorption (as in all panels of Fig. 19), that can contain emission blobs within the absorption area. This can be confused for real absorption in case there is a bright, warm background emission source such as a jet’s synchrotron in the background. Or it could be confused for a P-Cygni or inverse P-cygni profile. In our case, there are winds (Dasyra et al., 2015), but there is no detected synchrotron emission in the regions where the absorption is seen. (Note that the images and spectra shown in Fig. 19 are continuum subtracted.) Therefore, the absorption would have to be tracing outflows of cold, yet highly ionized gas that is absorbing the emission of hot dust in the background - a rather unusual configuration. However, we report the presence of the artifact for demonstration purposes.
B.2 Data with signal potentially linked to the cruciform artifact